OpenStudy (anonymous):

HELP If f(x) = 5x^2 − 4 and g(x) = x − 2, what is (f o g)(-3)?

OpenStudy (anonymous):

f o g is the same thing as saying f(g(x)) which looks a lot worse than it actually is. essentially this just means evaluate the inner function, then plug that result into the outer function. in this case the inner function is g(x) and you want to evaluate it at -3 g(-3) = x-2 = -3-2 = -5 then plug this value of g into f(g) to f(x) = 5x^2 - 4 = 5*(-5)^2 -4 = 5*25-4 = 125 - 4 = 121

OpenStudy (gorv):

fog(-3)=f(g(-3)) g(-3)=-3-2=-5 f(g(-3))=f(-5)=5*(-5)^2-4=5*25-4=125-4=121

OpenStudy (anonymous):

tHANKS SO MUCH GUYS